At equilibrium the concentrations in moles/L of N2, H2, and NH3 at 300 C are 0.25, 0.15, and 0.90 respectively. Find K for the reaction N2 + H2 ----- NH3


DEADLINE : 05/13/2022 11 : 00 PM




1
Expert's answer
2022-05-13T10:18:03-0400

"N_2+3H_2= 2NH_3"

"K=\frac{[NH_3]^2}{[N_2][H_2]^3}=\frac{0.90^2}{0.25 \times 0.15^3} = 0.011 L^2 mol^{-2}"